# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ II-C4 Interaction Module
The interaction module enables the LM agent to interact with humans, other agents, and the environment [41]. Through these varied interactions, the agent can gather diverse experiences and knowledge, which are essential for comprehensive understanding and adaptation.
  • Agent-Agent Interactions. The interaction module allows LM agents to communicate and collaborate with other agents, fostering a cooperative network where information and resources are shared [62]. This interaction can include coordinating efforts on shared tasks, exchanging knowledge to solve problems, and negotiating roles in multi-agent scenarios.
  • Agent-Human Interactions. LM agents can interact with humans including understanding and responding to natural language commands, recognizing and interpreting human emotions and expressions, and providing assistance in various tasks [20]. As observed, LLMs such as GPT-4 often tend to forget character settings in multi-turn dialogues and struggle with detailed role assignments due to context window limits. To address this, a tree-structured persona model is introduced in [65] for character assignment, detection, and maintenance, enhancing agent interactions.
@@ II-C5 Security Module
The security module is crucial to ensure the secure, safe, ethical, and privacy-preserving operations of LM agents [42]. It is designed to monitor and regulate the LM agent’s actions, interactions, and decisions to prevent harm and ensure compliance with legal and ethical standards. This module employs technologies such as hallucination mitigation, anomaly detection, and access control to identify and mitigate potential security/privacy threats. It also incorporates ethical guidelines and bias mitigation techniques to ensure fair and responsible behaviors. The security module can dynamically adapt to emerging threats by learning from new security/privacy incidents and integrating updates from security/privacy databases and policies.
@@
Connections Between Modules: The key components of an LM agent are interconnected to create a cohesive and intelligent system. Particularly, the planning module relies on the memory module to access past experiences and external knowledge, ensuring informed decision-making. The action module executes plans generated by the planning module, adapting actions based on real-time feedback and memory. The interaction module enhances these processes by facilitating communication and collaboration, which provides additional data and context for the planning and memory modules. Besides, security considerations are seamlessly integrated into every aspect of the LM agent’s operations to ensure robust and trustworthy performance.
Figure 8: Illustration of five underlying technologies, including their roles and key components, acting as the engine of the connected LM agents.
II-D Enabling Technologies of LM Agents
As illustrated in Fig. 8, there are five enabling technologies underlying the engine of connected LM agents.
Stage of LLM Description Utilized Technology Representive
Pre-training
Pre-train LLM in a self-supervised
manner on a large corpus
Transformer
GPT-3, PaLM-2,
LLaMA-2
Fine-tuning
Fine-tuning pre-trained
LLM for downstream tasks
Instruction-tuning,
Alignment-tuning,
Transfer learning
WebGPT[16], T0,
LLaMA-2-Chat
Prompting
Setup prompts and query trained
LLMs for generating responses
Prompt engineering,
Zero-shot prompting,
In-context learning
RALM[67],
ToolkenGPT[68]
II-D1 Large Foundation Model Technologies
LMs, such as LLMs and LVMs, serve as the core brains or controllers, providing advanced capabilities for AI agents across diverse applications [6, 57]. Table IV summarizes the basic training stages of LLM. (i) Multimodal capability: By employing multimodal perception (e.g., CLIP [69]) and tool utilization strategies, LM agents can perceive and process various data types from virtual and real environments, as well as human inputs, through multimodal data fusion, allowing for a holistic understanding of their surroundings [1]. Additionally, due to their powerful multimodal comprehension and generation capabilities [70], LM agents can interact seamlessly with both other agents and humans, fostering collaboration and competition among agents and humans. (ii) Advanced reasoning: LM technologies also empower AI agents with advanced reasoning and problem decomposition abilities [17]. For example, CoT [11] and ToT [12] reasoning allow LM agents to break down complex tasks into manageable sub-tasks, ensuring systematic and logical problem-solving approaches. (iii) Few/zero-shot generalization: Pre-trained on extensive corpora, LMs demonstrate few-shot and zero-shot generalization capabilities [25]. This allows LM agents to transfer knowledge seamlessly between tasks. (iv) Adaptability: Through continuous learning and adaptation, LM agents can accumulate knowledge and improve their performance over time by learning from new data and experiences [71]. This adaptability ensures that AI agents maintain high levels of performance in rapidly changing environments.
II-D2 Knowledge-Related Technologies
Knowledge-related technologies significantly enhance the capabilities of LM agents in knowledge representation [72], acquisition [73], fusion [74], retrieval [58], and synchronization [75]. These technologies collectively empower LM agents with deeper contextual understanding, more accurate reasoning, and better interaction. (i) Knowledge fusion: Knowledge representation and fusion technologies enable LM agents to build comprehensive knowledge bases by integrating information from diverse sources [74], such as structured databases, vector databases, and multimodal KGs. KGs, in particular, organize information into structured relationships, allowing LM agents to comprehend complex contexts and infer new information. (ii) RAG: RAG combines retrieval mechanisms with generative models, enabling LM agents to dynamically fetch relevant external information from vast knowledge sources [76]. This ensures that the outputs of LM agents are contextually relevant and up-to-date. Semantic understanding techniques [77, 78] further enhance LM agents’ ability to interpret and generate content that is both contextually appropriate and semantically rich. (iii) Knowledge synchronization: Continuous learning algorithms [79, 80] facilitate the ongoing updating of knowledge bases, allowing agents to incorporate new information incrementally and adapt to changing environments. These technologies also promote knowledge sharing and collaboration among multiple agents and between agents and humans.
II-D3 Interaction Technologies
Interaction technologies significantly enhance LM agents’ ability to engage with users, other agents, and the environment in natural, immersive, and contextually aware ways [42]. (i) HMI and human-robot interaction (HRI): HMI and HRI technologies [21], bolstered by NLP, enable LM agents to understand complex instructions, recognize speech, and interpret emotions, facilitating more intuitive interactions. (ii) 3D digital humans: 3D digital humans create realistic and engaging interfaces for communication, allowing LM agents to provide empathetic and personable interactions in applications such as customer support and healthcare [40]. (iii) AR/VR/MR: AR, VR, and MR technologies create immersive and interactive environments, allowing users to engage with digital and physical elements seamlessly. This is particularly beneficial in education, retail, and training, where such immersive experiences can enhance learning, engagement, and decision-making [48]. AR overlays digital information onto the real world, enabling users to interact with virtual objects in their physical surroundings; VR creates fully immersive digital environments for engaging with virtual elements; while MR blends digital information with physical environments. (iv) Multimodal interfaces: Multimodal interfaces enable LM agents to process and respond to various forms of input, including text, speech, images, gestures, and touch [1]. This allows users to interact with LM agents through their preferred modalities, making interactions more flexible and user-friendly.
II-D4 Digital Twin Technologies
Digital twin technologies [49] create comprehensive virtual representations of LM agents’ physical bodies and operational environments, continuously updated with real-time data from sensors, actuators, and other inputs. This ensures an accurate and up-to-date reflection of their real-world counterparts. (i) Virtual-physical synchronization: Digital twin technologies empower LM agents by enabling seamless and efficient synchronization of attributes, behaviors, states, and other data between their physical bodies and digital brains. This synchronization is achieved through intra-agent bidirectional communications, where the physical body continuously transmits real-time data to the digital twin for processing and analysis, while the digital twin sends back instructions and optimizations [49]. (ii) Virtual-physical feedback: This continuous feedback loop enhances LM agent’s contextual awareness, allowing for immediate adjustments and optimizations in response to changing conditions [23]. For example, an LM agent operating machinery can use its digital twin to anticipate mechanical wear and proactively schedule maintenance, thereby minimizing downtime and enhancing efficiency. (iii) Predictive analytics: Digital twins facilitate predictive analytics and simulation, enabling LM agents to anticipate future states and optimize their actions accordingly [22]. This capability is crucial in complex environments where unforeseen changes can significantly impact performance. Overall, digital twin technologies ensure that LM agents operate with high accuracy, adaptability, and responsiveness across diverse applications, effectively bridging the gap between the physical and digital realms.
II-D5 Multi-Agent Collaboration Technologies
Multi-agent collaboration technologies [41] enable coordinated efforts of multiple LM agents, allowing them to work together synergistically to achieve common goals and tackle complex tasks that would be challenging for individual agents to handle alone. (i) Data cooperation: It facilitates real-time and seamless information sharing through inter-agent communications, enabling LM agents to continuously synchronize their understanding of dynamic environments [81, 82]. (ii) Knowledge cooperation: By leveraging knowledge representation frameworks such as KGs [47, 83] and vector databases [84], LM agents can share and aggregate domain-specific insights, enhancing collective learning and decision-making. This shared knowledge base allows LM agents to build upon each other’s experiences to accelerate the learning process [71]. (iii) Computation cooperation: On the one hand, collaborative problem-solving techniques , such as multi-agent planning [85] and distributed reasoning [86], empower LM agents to jointly analyze complex issues, devise solutions, and execute coordinated actions. On the other hand, dynamic resource allocation mechanisms [87, 88], including market-based and game-theoretical approaches, enable LM agents to negotiate and allocate resources dynamically, thereby optimizing resource utilization ensuring effective task execution across multiple agents. This synergy is particularly beneficial in dynamic environments, which not only improves the operational capabilities of individual LM agents but also enhances the overall functionality of the system of connected LM agents.
TABLE V: Comparison of Existing Typical LM Agent Prototypes
@@ II-D5 Multi-Agent Collaboration Technologies
MobileAgent v2 Mobile communications Yes N/A Yes Multimodal input support
AppAgent Mobile communications Yes N/A Yes Performs tasks on mobile devices
Figure 02 by FigureAI Intelligent robots Yes N/A N/A Performs dangerous jobs
Optimus by Tesla Intelligent robots Yes N/A N/A Second-generation humanoid robot
@@
II-E Modern Prototypes & Applications of LM Agents
@@
Recently, various industrial projects of LM agents, such as AutoGPT, AutoGen, BabyAGI, ChatDev, and MetaGPT, demonstrate their diverse potential in assisting web, life, and business scenarios, such as planning personalized travels, automating creative content generation, and enhancing software development workflows. For example, AutoGPT [^3] https://autogpt.net/ is an open-source autonomous agent utilizing GPT-3.5 or GPT-4 APIs to independently execute complex tasks by breaking down them into several sub-tasks and chaining LLM outputs, showcasing advanced reasoning capabilities [14]. AutoGen [^4] https://microsoft.github.io/autogen/, developed by Microsoft, offers an open-source multi-agent conversation framework, supports APIs as tools for improved LLM inference, and emphasizes the automatic generation and fine-tuning of AI models [89]. BabyAGI [^5] https://github.com/yoheinakajima/babyagi integrates task management via OpenAI platforms and vector databases, simulating a simplified AGI by autonomously creating and executing tasks based on high-level objectives. ChatDev [^6] https://github.com/OpenBMB/ChatDev focuses on enhancing conversational AI, providing sophisticated dialogue management, coding, debugging, and project management to streamline software development processes [90]. MetaGPT [^7] https://www.deepwisdom.ai/ explores the meta-learning paradigm, where the model is trained to rapidly adapt to new tasks by leveraging knowledge from related tasks, thus improving efficiency and performance across diverse applications [91].
II-E1 Mobile Communications
LM agents offer significant advantages for mobile communications by enabling low-cost and context-aware decision-making [92], personalized user experiences [87], and automatic optimization problem formulation for wireless resource allocation [93]. For example, NetLLM [92] fine-tuning the LLM to acquire domain knowledge from multimodal data in networking scenarios (e.g., adaptive bitrate streaming, viewport prediction, and cluster job scheduling ) with reduced handcraft costs. Meanwhile, NetGPT [87] design a cloud-edge cooperative LM framework for personalized outputs and enhanced prompt responses in mobile communications via de-duplication and prompt enhancement technologies. ChatNet [94] uses four GPT-4 models to serve as analyzer (to plan network capacity and designate tools), planner (to decouple network tasks), calculator (to compute and optimize the cost), and executor (to produce customized network capacity solutions) via prompt engineering.
@@ II-E1 Mobile Communications
LM agents can also help enhance the QoE of end users. For example, MobileAgent v2 [18], launched by Alibaba, is a mobile device operation assistant that achieves effective navigation through multi-agent collaboration, automatically performing tasks such as application installation and map navigation, and supports multimodal input including visual perception, enhancing operational efficiency on mobile devices. AppAgent [19], developed by Tencent, performs various tasks on mobile phones through autonomous learning and imitating human click and swipe gestures, including posting on social media, helping users write and send emails, using maps, online shopping, and even complex image editing.
II-E2 Intelligent Robots
@@
LM agents play a crucial role in advancing intelligent industrial and service robots [21]. These robots can perform complex tasks such as product assembly, environmental cleaning, and customer service, by perceiving surroundings and learning necessary skills through deep learning models. In August 2024, FigureAI released Figure 02888https://spectrum.ieee.org/figure-new-humanoid-robot, a human-like robot powered by OpenAI LM, capable of fast common-sense visual reasoning and speech-to-speech conversation with humans to handle dangerous jobs in various environments. Besides, in July 2024, Tesla unveiled its second-generation humanoid robot named Optimus [^9] https://viso.ai/edge-ai/tesla-bot-optimus/, which demonstrates the enhanced capabilities brought by advanced LM agents.
II-E3 Autonomous Driving
@@
LM agents are transforming autonomous driving by enhancing vehicle intelligence, improving safety, and optimizing driving experience (e.g., offering personalized in-car experience) [22]. In May 2024, Baidu’s Apollo Day saw the launch of Carrot Run’s sixth-generation unmanned vehicle, which is built upon the Apollo Autonomous Driving Foundation Model (ADFM) [^10] https://autonews.gasgoo.com/icv/70033042.html, the first LM agent supporting L4 autonomous driving. Companies such as Tesla, Waymo, and Cruise are also leveraging LM agents to refine their autonomous driving systems, aiming for safer and more efficient transportation solutions.
II-E4 Autonomous attack-defense confrontation

### v2

@@ II-C4 Interaction Module
The interaction module enables the LM agent to interact with humans, other agents, and the environment [44]. Through these varied interactions, LM agents can gather diverse experiences and knowledge, which are essential for comprehensive understanding and adaptation.
  • Agent-Agent Interactions. LM agents can coordinate efforts on shared tasks, exchange knowledge to solve problems, and negotiate roles to collaborate with other agents, fostering a cooperative network [63]. Theory of mind [64] can enhance the ability to understand other agents’ hidden mental states. Li et al. [64] propose a prompt-engineering method to incorporate explicit belief state representations. They also introduce a novel evaluation of LLMs’ high-order theory of mind in teamwork scenarios, emphasizing dynamic belief state evolution and intent communication among agents.
  • Agent-Human Interactions. LM agents can interact with humans by understanding natural languages, recognizing human emotions and expressions, and providing assistance in various tasks [20]. As found by work [65], LLMs such as GPT-4 tend to forget character settings in multi-turn dialogues and struggle with detailed role assignments due to context window limits. To address this, a tree-structured persona model is devised in [65] for character assignment, detection, and maintenance, enhancing agent interactions.
@@ II-C5 Security Module
The security module is crucial to ensure the secure, safe, ethical, and privacy-preserving operations of LM agents [45]. It is designed to monitor and regulate the LM agents’ actions, interactions, and decisions to prevent harm and ensure compliance with legal and ethical standards. LMs can be utilized to design advanced and cost-effective defense mechanisms, such as PentestGPT [22] for automated network vulnerability detection. LMs also introduce new threats to agents such as magnified hallucination and memorization risks. The security module should dynamically adapt to emerging threats by learning from new security/privacy incidents and integrating updates from security/privacy databases and policies. It also incorporates ethical guidelines and bias mitigation techniques to ensure fair and responsible behaviors.
@@
Connections Between Modules: The above five key components of an LM agent are interconnected to create a cohesive and intelligent system. Particularly, the planning module relies on the memory module to access past experiences and external knowledge, ensuring informed decision-making. The action module executes plans generated by the planning module, adapting actions based on real-time feedback and memory. The interaction module enhances these processes by facilitating communication and collaboration, which provides additional data and context for the planning and memory modules. Besides, security considerations are seamlessly integrated into every aspect of the LM agent’s operations to ensure robust and trusted executions.
TABLE III: Comparison of Existing Typical LM Agent Prototypes
@@
MobileAgent v2 General scenarios Yes N/A Yes Multimodal input support
AppAgent General scenarios Yes N/A Yes Performs tasks on mobile devices
@@
Figure 02 by FigureAI Intelligent robots Yes Yes N/A Performs dangerous jobs
Optimus by Tesla Intelligent robots Yes Yes N/A Second-generation humanoid robot
@@
II-D Modern Prototypes & Applications of LM Agents
II-D1 General Scenarios
@@
Recently, various industrial projects of LM agents, such as AutoGPT, AutoGen, BabyAGI, ChatDev, and MetaGPT, demonstrate their diverse potential in assisting web, life, and business scenarios, such as planning personalized travels, automating creative content generation, and enhancing software development workflows. For instance, AutoGPT [^3] https://autogpt.net/ is an open-source autonomous agent utilizing GPT-3.5 or GPT-4 APIs to independently execute complex tasks by breaking down them into several sub-tasks and chaining LLM outputs, showcasing advanced reasoning capabilities [12]. AutoGen [^4] https://microsoft.github.io/autogen/, developed by Microsoft, offers an open-source multi-agent conversation framework, supports APIs as tools for improved LLM inference, and emphasizes the automatic generation and fine-tuning of AI models [13]. BabyAGI [^5] https://github.com/yoheinakajima/babyagi integrates task management via OpenAI platforms and vector databases, simulating a simplified AGI by autonomously creating and executing tasks based on high-level objectives. ChatDev [^6] https://github.com/OpenBMB/ChatDev focuses on enhancing conversational AI, providing sophisticated dialogue management, coding, debugging, and project management to streamline software development processes [67]. MetaGPT [^7] https://www.deepwisdom.ai/ explores the meta-learning paradigm, where the model is trained to rapidly adapt to new tasks by leveraging knowledge from related tasks, thus improving performance across diverse applications [68].
@@
LM agents can also enhance the QoE of end users. For instance, MobileAgent v2 [18], launched by Alibaba, is a mobile device operation assistant that achieves effective navigation through multi-agent collaboration, automatically performing tasks such as map navigation, and supports multimodal input including visual perception, enhancing operational efficiency on mobile devices. AppAgent [19], developed by Tencent, performs various tasks on mobile phones through autonomous learning and imitating human click and swipe gestures, including posting on social media, helping users write and send emails, using maps, online shopping, and even complex image editing.
II-D2 Mobile Communications
LM agents offer significant advantages for mobile communications by enabling low-cost and context-aware decision-making [69], personalized user experiences [70], and automatic optimization problem formulation for wireless resource allocation [71]. For instance, NetLLM [69] fine-tunes the LLM to acquire domain knowledge from multimodal data in networking scenarios (e.g., adaptive bitrate streaming, viewport prediction, and cluster job scheduling) with reduced handcraft costs. Meanwhile, NetGPT [70] uses a cloud-edge cooperative LM framework for personalized outputs and enhanced prompt responses in mobile communications via de-duplication and prompt enhancement technologies. ChatNet [72] uses four GPT-4 models to serve as analyzer (to plan network capacity and designate tools), planner (to decouple network tasks), calculator (to compute and optimize the cost), and executor (to produce customized network capacity solutions) via prompt engineering.
II-D3 Intelligent Robots
@@
LM agents play a crucial role in advancing intelligent industrial and service robots [15]. These robots can perform complex tasks such as product assembly, environmental cleaning, and customer service, by perceiving surroundings and learning necessary skills through LMs. In August 2024, FigureAI released Figure 02888https://spectrum.ieee.org/figure-new-humanoid-robot, a human-like robot powered by OpenAI LM, capable of fast common-sense visual reasoning and speech-to-speech conversation with humans to handle dangerous jobs in various environments. Besides, in July 2024, Tesla unveiled its second-generation humanoid robot named Optimus [^9] https://viso.ai/edge-ai/tesla-bot-optimus/, which demonstrates the enhanced capabilities brought by LM agents.
II-D4 Autonomous Driving
@@
LM agents are transforming autonomous driving by enhancing vehicle intelligence, improving safety, and optimizing driving experience (e.g., offering personalized in-car experience) [21]. In May 2024, Baidu’s Apollo Day saw the launch of Carrot Run’s sixth-generation unmanned vehicle, which is built upon the Apollo autonomous driving foundation model (ADFM) [^10] https://autonews.gasgoo.com/icv/70033042.html, the first LM agent supporting L4 autonomous driving. Companies such as Tesla, Waymo, and Cruise are also leveraging LM agents to refine their autonomous driving systems, aiming for safer and more efficient transportation systems.
Figure 11: General architecture of LM agent networks in bridging the human, physical, and cyber worlds. The LM agent has five constructing modules: planning, action, memory, interaction, and security modules. The engine of connected LM agents is empowered by a combination of five cutting-edge technologies: foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration. LM agents interact with humans through HMI technologies such as NLP, with the assistance of handheld devices and wearables, to understand human intentions, desires, and beliefs. LM agents can synchronize data and statuses between the physical body and the digital brain through digital twin technologies, as well as perceive and act upon the surrounding virtual/real environment. LM agents can be interconnected in the cyber space through efficient cloud-edge networking for efficient sharing of data, knowledge, and computation results to promote multi-agent collaboration.
II-D5 Autonomous attack-defense confrontation